OpenStudy (iwanttogotostanford):

PLEASE HELP URGENT: For the function f(x) = 3(x − 1)2 + 2, identify the vertex, domain, and range.

OpenStudy (iwanttogotostanford):

@mathway @myininaya @CloverRacer @l

OpenStudy (iwanttogotostanford):

@shelby1290 @Bee_see @wolf1728

OpenStudy (iwanttogotostanford):

@vera_ewing

OpenStudy (wolf1728):

Would that be: f(x) = 3*(x − 1)*(x − 1) + 2 ?

OpenStudy (iwanttogotostanford):

yes

OpenStudy (iwanttogotostanford):

well it i this; f(x) = 3(x − 1)^2 + 2 sorry

OpenStudy (sleepyhead314):

The vertex form of a parabola is given as \( y = a (x-h)^2 + k \) where the point \( (h,k) \) is the vertex so comparing the two: \(y=3(x-1)^2 +2\) and \(y=a(x-h)^2+k\) can you tell me what you think the vertex is? :)

OpenStudy (iwanttogotostanford):

the first one?

OpenStudy (wolf1728):

Hmm, it seems sleepyhead is trying another approach, I multiplied it out: f(x) = 3*(x − 1)*(x − 1) + 2 equals 3 * x^2 -2x +1 +2 equals 3 x^2 -2x +3

OpenStudy (sleepyhead314):

I mean as a point :P or tell me h = k = that's ok too

OpenStudy (sleepyhead314):

yes, as it asks for the vertex - I assumed it would be easier to point it out since it is already in "vertex form" @wolf1728 :)

OpenStudy (iwanttogotostanford):

these are my choices: The vertex is (1, 2), the domain is all real numbers, and the range is y ≥ 2. The vertex is (1, 2), the domain is all real numbers, and the range is y ≤ 2. The vertex is (–1, 2), the domain is all real numbers, and the range is y ≥ 2. The vertex is (–1, 2), the domain is all real numbers, and the range is y ≤ 2.

OpenStudy (sleepyhead314):

then yes, your guess is correct! @iwanttogotostanford Good Job! :)

OpenStudy (iwanttogotostanford):

so it would be A?:_)

OpenStudy (sleepyhead314):

indeed ^_^

OpenStudy (iwanttogotostanford):

WAIT I HAVE ANOTHER QUSTION DONT LEAVE

OpenStudy (iwanttogotostanford):

thank you WAIT CAN I ASK ANOTHER PLEASE

OpenStudy (iwanttogotostanford):

Given the functions f(x) = 3x2, g(x) = x2 − 4x + 5, and h(x) = –2x2 + 4x + 1, rank them from least to greatest based on their axis of symmetry.

OpenStudy (iwanttogotostanford):

these are my choices:-) thanks for the help f(x), g(x), h(x) f(x), h(x), g(x) g(x), h(x), f(x) g(x), f(x), h(x)

OpenStudy (sleepyhead314):

the axis of symmetry is given by \( \large x = \frac{-b}{2a}\) when the equations are in a form \(y=ax^2+bx+c\)

OpenStudy (iwanttogotostanford):

so it would be c? @sleepyhead314

OpenStudy (sleepyhead314):

yes :) that's what I got too! you're doing great ^_^

OpenStudy (iwanttogotostanford):

thank you! one more?? sorry.... @wolf1728 @sleepyhead314

OpenStudy (iwanttogotostanford):

Given the function f(x) = x2 and k = –1, which of the following represents a function opening downward? Answer choices: f(x) + k kf(x) f(x + k) f(kx)

OpenStudy (sleepyhead314):

perhaps consider making a new question so that we don't both wolf too much with the notifications :)

OpenStudy (iwanttogotostanford):

okay i will!

OpenStudy (wolf1728):

I made a mistake on previous problem f(x) = 3*(x − 1)*(x − 1) + 2 equals 3 * (x^2 -2x +1) +2 equals 3 x^2 -6x +5